OpenStudy (anonymous):

(3r^4-2r^3+2r^2+12r-6)+(r^5+6r^3+7r^2+12r-6)+(r^5+6r^3+7r^2-2r+4)+(-6r^4+r^2-8r-10)

OpenStudy (anonymous):

no it doesn't

OpenStudy (dumbcow):

"no it doesn't" lol

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

Im glad that you all find me amusing....but I can do with without the laughter.

OpenStudy (saifoo.khan):

We just have to add like terms, x^4 with x^4, x^3 with x^3 and so on, Therefore, \[r^ 5 - 3r^4 +10r^3 + 17r^2+14r -1 8\]

OpenStudy (anonymous):

and with all the laughter, the answer key indicates that you are wrong.

OpenStudy (saifoo.khan):

are u sure the question you typed is correct? o_O

OpenStudy (anonymous):

of course

OpenStudy (saifoo.khan):

i can bet im correct.

OpenStudy (dumbcow):

you guys missed a " r^5 " \[\rightarrow 2r^{5}-3r^{4}+10r^{3}+17r^{2}+14r-18\] @mslorraine no disrespect intended, the bluntness in which you said it made me think of something else

OpenStudy (anonymous):

find the variation constant and equation of variation where y varies directly as x and y=20 when x=20

OpenStudy (dumbcow):

y= kx --> k = y/x = 20/20 = 1 y = x